want to check my answer on....
3t 1
__ = __
10 2
I said it was t = 0.6

Answer by checkley77(12844)

3t 1
__ = __
10 2
3t*2=10*1
6t=10
t=10/6=5/3=1.6667 answer.
Proof:
3*1.6667/10=1/2
5/10=1/2
1/2=1/2